OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary, round answers to the nearest hundredth. 5x2 + 16x – 84 = 0

Nnesha (nnesha):

\[\huge\ x =\frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\] this is quadratic formula

Nnesha (nnesha):

\[Ax^2+Bx+C=0\]this is quadratic equation now replace a b and c by their values that you have inthe equation

Nnesha (nnesha):

 now replace